OpenStudy (anonymous):

What is the solution to the equation log[2] 4 - log[2] x = 3 ?

OpenStudy (anonymous):

write it as one log_2 : log_2 (4/x) = 3 now write this in exponential form: 2^3 = 4/x can you solve from here?

OpenStudy (anonymous):

hmm is it 32?

OpenStudy (anonymous):

no... cross multiply to get a new equation: x = 4/(2^3)

OpenStudy (anonymous):

no x=1/2

OpenStudy (anonymous):

Oh ok thanks:)